the problem is 4x-6y=5and 2x-3y=7 am I suppose to multiple the top by a number

Answer by richwmiller(17219)

4x-6y=5and 2x-3y=7
You haven't told us what you are trying to do.
4x-6y=5
2x-3y=7
But you may notice that they have the same slope and so they are parallel and never intersect and therefore have no "solution"
-a/b= slope
-4/-6=2/3
-2/-3=2/3